the diffrence between six times a number and 9 is equal to five times the sum of the number and 2. find the number: the diffrence between six times a number and 9 is equal to five times the sum of the number and 2. find the number
Answer by orca(336)

Let x be the number.
Then
The difference between six times a number and 9 can be expressed as
6x - 9
Five times the sum of the number and 2 can be expressed as
5(x + 2)
As they are the same, we have
6x - 9 = 5(x + 2)
Solving for x, we have
6x - 9 = 5x + 10
x = 19
So the number is 19.